OpenStudy (anonymous):

solve for x: ln(-5x+3)=ln2x+2

OpenStudy (kropot72):

ln(-5x + 3) = ln(2x) + 2 Subtracting ln(2x) from both sides gives ln(-5x + 3) - ln(2x) = 2 By the rules of logarithms \[\ln \frac{-5x+3}{2x}=2\] Therefore \[e ^{2}=\frac{-5x+3}{2x}\] \[(7.389\times 2)x=-5x+3\] Now you just need to solve for x.